OpenStudy (anonymous):

can any one help me in this Exercises Find any horizontal and vertical asymptotes and any holes that may exist for each rational function y = 2 / 3+2x

OpenStudy (anonymous):

you need the graph?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

3+2x=0 , \[ x \neq-\frac{3}{2}\] vertical asymptotes:x= -3/2 degree domain and number difference than no horizontal asymptotes , no holes \[x-intercept : 0= \frac{2}{3+2x}= 0*(3+2x) =2\] x -intercept = o or none Find : y - intercept\[y=\frac{2}{3+(2*0)}= \frac{2}{3}\]

OpenStudy (anonymous):

asymptotes are the x axis and x=-1.5

OpenStudy (anonymous):